For any z∈C, if |z| , then ∈R z∈ R . Is it true or false ? Give reasons.
1
Expert's answer
2020-02-27T11:29:59-0500

Let "z=a+bi \in \mathbb{C}," then "\left|z\right| = \sqrt{a^2+b^2}". Therefore, the absolute value of a complex number always lies in "\mathbb{R}."


Answer: False.